OpenStudy (anonymous):

Graph the system of inequalities y>3 x+1 y <= -2/3x+4 what are the two points? I keep getting (9,11)

OpenStudy (anonymous):

ok, first graph y=3x+1 y=-2/3 x + 4

OpenStudy (anonymous):

that doesnt help me with the points....

jimthompson5910 (jim_thompson5910):

To find points, plug in values for x to find corresponding values for y Ex: Plug in x = 0 in y=3x+1 to get y=3x+1 y=3(0)+1 y=0+1 y = 1 So when x = 0, we have y = 1. This gives us one point (0, 1) Now plug in x = 1 y=3x+1 y=3(1)+1 y=3+1 y=4 so another point is (1, 4) This tells us that the two points (0,3) and (1,4) lie on the line y = 3x+1

jimthompson5910 (jim_thompson5910):

You only need two points to create any line and you can plug in any value you want for x (it's best to keep x small so calculations are easy and you'll be able to see it on the graph)

OpenStudy (anonymous):

so any one of those would work for both the equations?

jimthompson5910 (jim_thompson5910):

no, both are necessary to graph y = 3x+1 you would follow the same steps to graph y=-2/3 x + 4

OpenStudy (anonymous):

but i need only one point for both of the equations.

jimthompson5910 (jim_thompson5910):

its true that both equations intersect at one point, but it's not essential to use that point to graph you can use any two points on a line to make that line (and it doesn't have to be the point of intersection)

OpenStudy (anonymous):

i dont understand what i would put for the answer then....becuz its also telling me to substitue the points to prove that they work for both equations

jimthompson5910 (jim_thompson5910):

so it's asking you to graph and to test points (that aren't on the graph)?

OpenStudy (anonymous):

here ill give you what it says. Graph the system of inequalities .Please provide only one ordered pair (x, y) that will work for this system of inequalities. prove your answer by substituting the (x, y) order pair into both inequalities.

jimthompson5910 (jim_thompson5910):

Ah ok, that makes more sense This problem is basically two problems thrown together. The first part has you graph the system. You can then use the graph to find a point that lies in the solution region. Once you've identified a point that lies in the solution region, you would then plug it into both inequalities to test that point.

jimthompson5910 (jim_thompson5910):

So have you graphed this system of inequalities?

OpenStudy (anonymous):

yea it gives me some werid point (9,11) which i dont think is right

jimthompson5910 (jim_thompson5910):

No, (9,11) is not in the solution region, so you're correct there. Have you identified or shaded the solution region?

OpenStudy (anonymous):

No id ont know how to do that...

jimthompson5910 (jim_thompson5910):

Do you know how to graph y = 3x+1?

OpenStudy (anonymous):

not really...prolly why i got it wrong

jimthompson5910 (jim_thompson5910):

look at my post above and I show you that plugging in x = 0 gives you y = 1. So we have one point (0,1) Similarly, we have another point (1,4) So plot these two points and draw a straight line through them to graph y=3x+1